OpenStudy (anonymous):

7x = 20 + 4x It says my answer options are : (A) 5 and 2/3 , (B) 6 and 1/3 , (C) 6 and 2/3 , and (D) 28?

OpenStudy (anonymous):

only one answer that is 20/3 7x=20+4x Since 4x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 4x from both sides. 7x-4x=20 Since 7x and -4x are like terms, add -4x to 7x to get 3x. 3x=20 Divide each term in the equation by 3. (3x)/(3)=(20)/(3) Simplify the left-hand side of the equation by canceling the common terms. x=(20)/(3)

OpenStudy (anonymous):

so it is c

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

6+2/3 = 20/3 right!

OpenStudy (anonymous):

thank you (: I feel hopeless. I'm taking a test and there's so many questions like this that I can't answer ):